### v1

@@ 4 Experiments
4.1 Experiment Setup
In experiment part, we mainly focus on models that either demonstrate advanced general capability or are explicitly pre-trained for code generation. We evaluate both close-sourced and open-sourced models. The Closed-source models include o1-mini, GPT-4o-mini, GPT-4o, and Claude-3.5-Sonnet and the open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct[9], Qwen-2.5-72B-Instruct[26] and QwQ-32B[24]), the DeepSeek family (DeepSeek-v3[18] and DeepSeek-R1[6]), the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct[4]) and Yi-Coder-9B-Chat[20]. Together, these models cover the most widely used large-parameter and reasoning models in code generation task.
To reduce evaluation overhead and eliminate potential data leakage, we select the most recent 20% of problems in CodeFlowBench as our test set. Each model is evaluated in both multi-turn and single-turn patterns, using Pass@k and APD as the primary metrics. Detailed experimental settings is included in appendix.
4.2 Main Result
Table 1: Performances of pass@1 and Average Pass Depth of various models on CodeFlowBench. The overall average depth of problems in CodeFlowBench is 1.94.
Model pass@1 Average Pass Depth
@@ 4.2 Main Result
GPT-4o 17.5 30.2 0.494 0.177 0.539 0.758 1.042
Claude-3.5-Sonnet 13.6 36 0.414 0.117 0.444 0.648 1.167
Open-Sources (7B-Level)
Qwen2.5-Coder-7B-Instruct 2.3 15 0.233 0.018 0.247 0.436 0.750
@@
Yi-Coder-9B-Chat 5 13.8 0.228 0.021 0.270 0.388 0.500
Open-Sources (32B-Level)
@@
Open-Sources (70B-Level)
Llama-3.3-70B-Instruct 15 27.6 0.448 0.163 0.465 0.733 1.042
@@
Open-Sources (Large Model)
Deepseek-V3 18 35.7 0.529 0.219 0.549 0.836 1.208
@@
Performances of different LLMs
Table 1 shows the main result of our experiment. The best pass@1 achieved on CodeFlowBench is only 20.8% and the highest Average Pass Depth (APD) is also only 0.541, underscoring the benchmark is challenging enough for today’s leading LLMs. It follows our intuition that closed-source models generally outperform open-source models.However, the gap is not decisive: the open-source DeepSeek-V3 exceeds most closed-source competitors, and Llama-3.3-70B-Instruct remains highly competitive. Besides, two kinds of model consistently punch above their parameter class: specialized coding models and reasoning models. Smaller coder models like Qwen2.5-Coder and Yi-Coder deliver notably strong pass@1 scores compared to large models, while reasoning models like QwQ-32B-Preview can approach the performance of GPT-4o. These results suggest that advancing domain-specific coding architectures and refining reasoning strategies are promising directions for future performance gains.
Comparison of pass@k on multi-turn and single-turn pattern
The results in Table 1 reveal a substantial gap between the multi-turn and single-turn settings, indicating that even when facing the same underlying problem, models find multi-turn cases significantly more challenging to handle. Furthermore, we observe distinct trends in model performance across the two patterns. Models such as Claude-3.5-Sonnet and Deepseek-R1 achieve outstanding results under the single-turn setting, yet their performance in multi-turn scenarios is less remarkable. In contrast, models like GPT-4o and Llama-3.3-70B-Instruct show relatively smaller performance gaps between the two settings. Unlike the single-turn pattern, which primarily requires complex reasoning, the multi-turn pattern demands that models generate reusable and generalizable functions based on a given dependency structure and pre-implemented materials, thus requiring more comprehensive capabilities. As a result, different models exhibit varying shortcomings that need to be further addressed and enhanced.
Comparison of APD on different depth
Results from APD provide additional valuable insights. We observe that small open-source models exhibit varied strengths between low-depth problems and deep-depth problems. Models like Qwen2.5-Coder and Llama-3 perform better on problems whose depth beyond 3, whereas Yi-Coder perform better on problems whose depth is 1 or 2. In contrast, large and closed-source models with higher overall APD demonstrate more balanced performance, typically passing 20–30% depth of the whole problems in both low-depth problems and deep-depth problems. Additionally, while reasoning model like Deepseek-R1 and o1-mini achieve SOTA performance on our benchmark, it doesn’t reflect obvious advantage on deep-depth problems, indicating a possible direction for reasoning model to make thinking chain fit multi-turn cases better.
5 Discussion
While models may succeed on multi-turn tasks with a small number of turns and simple linear dependencies structure, they are nearly incapable of solving all turns correctly when the structure of problem becomes more complex.
The results in Table 1 report the average pass@1 scores across models. However, a deeper look of the problems that were successfully solved reveals a clear imbalance: most of the passed cases correspond to problems with simple dependency structures. As the problem structure becomes more complex, existing SoTA models struggle significantly and are nearly impossible for them to correctly complete all turns and solve the entire problem.
Figure 5: The pass@1 score of models on multi-turn problems with different turns.We divide models into 4 groups: closed-source models, open-source 7B-level models, open-source 32B&70B-level models and open-source large models & reasoning models.
Figure 5 presents the pass@1 scores of models on multi-turn problems across different numbers of turns. All curves, regardless of model group, exhibit a consistent trend: they begin at relatively high scores for problems with only 1 or 2 turns, but drop sharply as the number of turns increases, eventually reaching zero. For small open-source models, the scores remain low across all turn depths. Although larger open-source models and closed-source models perform significantly better on short-turn problems, their performance still degrades notably on problems with more turns. Even top-performing models such as o1-mini and Deepseek-R1, which achieve strong overall pass@1 scores, fail to solve problems with more than six turns. This result highlights a key limitation: current models still struggle to generate code in each turn that generalizes well, covers a wide range of cases, and integrates seamlessly with higher-level functions.
To further measure the complexity of dependency structure, we introduce a metric called DSC(Dependency Structure Complexity) which is defined as the ratio of its overall turns to the number of overall depth:
$\mathrm{DSC}(\text{problem})=\frac{\mathrm{OverallTurn}(\text{problem})}{\mathrm{OverallDepth}(\text{problem})}$
Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST. Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure. Figure 6 presents the pass@1 scores of models across different DSC intervals. It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure. Only a few leading models, such as o1-mini, GPT-4o and Deepseek-V3, are able to solve a limited number of problems with DSC values below 1.33. All models struggle significantly when faced with problems involving more complex structures. These results highlight the current models’ limited ability to understand intricate dependency relationships and to generate functions that can effectively handle them.
Which Kind of Error Lead to LLMs’ Poor Performance in Multi-Turn Pattern?
Since a huge gap between models’ performances on multi-turn and single-turn pattern is demonstrated, we may be curious about the underlying reasons. Firstly, we define a taxonomy where errors are concluded into three types:
  • Incomplete Reasoning (IR) The model often handles only straightforward “happy-path” cases and fails to generalize. It may misinterpret or oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapse on larger inputs. This reflects limitation of model in its reasoning ability.
  • Insufficient Globalization (IG) The function’s logic may run correctly in isolation but omits necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This reflects limitation of model in managing global context.
  • Instruction Misinterpretation (IM) In multi-step or multi-turn prompts, the model may solve isolated subtasks but fail to respect overall orchestration: it can overlook how to invoke provided helper functions and arrange the codeflow properly in high-level function, leading to integration failures. This reflects limitation of model in task understanding and dependency relationship handling.
Compared to the single-turn pattern, Incomplete Reasoning (IR) remains a shared challenge where the single-turn approach can leverage holistic code coordination and design to address logical issues that localized implementations may fail to anticipate; Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are problems specific to the multi-turn pattern. We then conduct a study to quantify the distribution of each error type made by the LLMs. We randomly sample examples from the test set and manually annotate their error categories. The results are summarized in Table 2.
Table 2: Proportion of each error type across Deepseek-V3 and o1-mini.IR, IG and IM stand for Incomplete Reasoning, Insufficient Globalization and Instruction Misinterpretation, respectively.
Model IR(%) IG(%) IM(%) Others(%)
Deepseek-v3 47.2 14.6 32.6 5.6
o1-mini 38.7 11.8 43.0 6.5
From Table 2, we draw several insights. First, Incomplete Reasoning (IR) remains the dominant error type which is the main challenge in competition-level code generation. We need to distinguish the effect of multi-turn setting out by quantifying how much of IR can be “fixed” by switching from multi-turn pattern to single-turn pattern. Therefore, we introduce the IR Remediation Rate, which is defined as the fraction of IR errors under the multi-turn pattern that are correctly answered under single-turn evaluation. We report the IR remediation rate of two models tested above in Table 3.
Table 3: IR remediation rate of Deepseek-V3 and o1-mini.
Model IR Remediation Rate (%)
DeepSeek-v3 16.7
o1-mini 8.3
By contrast, Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are largely specific to multi-turn interactions. This makes sense: CodeFlowBench tasks require multi-step orchestration, where the model must (1) decompose the high-level problem into subtasks, (2) correctly invoke pre-implemented functions and respect their dependencies, and (3) assemble an end-to-end solution that can be executed iteratively. The high rate of IM errors highlights that models still struggle with consistent instruction following across multiple turns—understandable, given that they are typically trained on isolated, single-turn examples. Finally, we observe that models often lack “global awareness” when generating low-level code: they may omit necessary imports or mishandle shared state, even though tool functions and the top-level wrapper are expected to integrate seamlessly. This underlines the importance of evaluating code generation in settings that demand both local correctness and global coherence.
Appendix B presents a more detailed case study of these errors.
@@ 6 Conclusion
In this paper, we introduce CodeFlowBench, the first benchmark specifically designed for multi-turn, iterative code generation. Comprising 5,258 complex programming problems sourced from Codeforces, CodeFlowBench is underpinned by a novel automatic data curation pipeline and a comprehensive evaluation framework that rigorously assesses a model’s ability in multi-turn and iterative code generation. Our experiment across a range of popular LLMs highlights the substantial challenges posed by this benchmark: current leading models struggle markedly with multi-turn tasks, exhibiting a pronounced performance gap compared to single-turn generation. While different models demonstrate strengths on problems with different depth, almost none can fully and correctly resolve structurally complex tasks once. Besides, we analyze key error types contributing to these shortcomings and discuss their implications. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also provides a valuable foundation for future research aimed at closing this gap and supporting realistic development workflows.
Limitation & Future Work
While inspired by real-world development workflows, our current dataset remains confined to competition-level problems at function and class level. In future work, we plan to expand CodeFlowBench to repository-level challenges, which feature far more intricate dependency structures and demand explicit codeflow organization to ensure extensibility and maintainability. Additionally, we intend to design and train a dedicated code-generation model specifically tailored to this multi-turn, dependency-aware setting.
Appendix A Data Curation Pipeline
A.1 Detail of Problem Scraping
@@
A.2 Detail of Solution Scraping
@@
A.3 Detail of Solution Code Generation
@@
A.4 Detail of Subproblems Generation
@@
A example of subproblems in shown in Figure 13
@@
A.5 Detail of Test Cases Generation
@@
Appendix B Error Case Study

### v2

@@ 4 Experiments
Experiment Setup
For comprehensive experiments, we evaluate both close-sourced and open-sourced models. The closed-source models include the GPT family (o3-mini, o1-mini, GPT-4.1-mini, GPT-4o-mini and GPT-4o) [19, 17], Claude-3.5-Sonnet and Gemini-2.0-flash. The open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct, Qwen-2.5-72B-Instruct and QwQ-32B) [16, 37, 34], the DeepSeek family (DeepSeek-v3 and DeepSeek-R1) [27, 11], the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct) [8] and Yi-Coder-9B-Chat [30]. To eliminate potential data leakage and unnecessary evaluation overhead, we select the most recent 1000 problems from CodeFlowBench for tests. Each model is evaluated in both multi-turn and single-turn scenarios, with Pass@k and APD@k as the primary metrics. Please refer to Appendix D.1 and D.2 for more implementation details.
4.1 Main Experiments
Table 1: Performance comparison on CodeFlowBench with Pass@1 and APD@1 (Average Pass Depth) metrics. As a reference for overall APD, CodeFlowBench’s average problem depth is 1.94.
Model Pass@1 APD@1 (Average Pass Depth)
@@ 4.1 Main Experiments
Gemini-2.0-flash 17.3 31.6 0.496 0.183 0.527 0.782 1.042
Claude-3.5-Sonnet 13.6 36.0 0.414 0.117 0.444 0.648 1.167
Open-Source (7B-Level)
@@
Qwen2.5-Coder-7B-Instruct 2.3 15.0 0.233 0.018 0.247 0.436 0.750
Yi-Coder-9B-Chat 5.0 13.8 0.228 0.021 0.270 0.388 0.500
Open-Source (32B-Level)
@@
Open-Source (70B-Level)
Llama-3.3-70B-Instruct 15.0 27.6 0.448 0.163 0.465 0.733 1.042
@@
Open-Source (Large Model)
Deepseek-V3 18.0 35.7 0.529 0.219 0.549 0.836 1.208
@@
Overall Performance
Table 1 presents the extensive experiments across 1,000 latest problems. The highest observed Pass@1 score of 24.4% and maximum APD@1 of 0.602 demonstrate the benchmark’s rigorousness, with even state-of-the-art LLMs struggling to achieve satisfactory performance. While closed-source models generally outperform the open-source ones, exceptions exist—notably DeepSeek-R1 surpasses several proprietary models. Notably, two model categories exhibit parameter-efficient competence: specialized coding models and reasoning models. Small coding models like Qwen2.5-Coder and Yi-Coder deliver surprising Pass@1 relative to model size, while reasoning models like QwQ-32B-Preview approach GPT-4o’s performance via test-time scaling. These findings suggest that domain-specific and reasoning models are promising directions for future advancements.
Multi-Turn versus Single-Turn
Table 1 also reveals the substantial performance gap between the multi-turn and single-turn scenarios. This gap underscores the inherent complexity of multi-turn code generation, which requires models to "look before and after" for long-context coherence across iterative function implementations. Furthermore, we observe distinct features of models’ performance across single-turn and multi-turn scenarios: while models like Claude-3.5-Sonnet and DeepSeek-R1 excel in single-turn scenarios, they exhibit up to +60% performance degradation in multi-turn settings. In contrast, models such as GPT-4o and Llama-3.3-70B-Instruct demonstrate more consistent performance. The fundamental difference lies in the cognitive demands: single-turn evaluation examines complex reasoning capabilities, while multi-turn evaluation needs to generate reusable, modular functions, and integrate them into larger components. These insights expose under-explored limitations such as iterative development and dependency management.
Table 2: Performance comparison on CodeFlowBench with the APT@1 (Average Pass Turn) metric. As a reference for overall APT, CodeFlowBench’s average problem turn is 2.20.
Model Overall Turn1 Turn2 Turn3 Turn4 Turn5
Closed-Source
o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
Open-Source (7B-Level)
Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
Open-Source (32B-Level)
Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
Open-Source (70B-Level)
Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
Open-Source (Large Model)
Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
Depth-Wise Performance
The performance stratification across dependency depths reveals distinct model characteristics. Small open-source models (7B-32B parameters) show pronounced depth-specific biases—Qwen2.5-Coder-7B achieves 0.750 APD@1 on Depth4 problems versus 0.018 on Depth1, suggesting emergent compositional abilities despite limited parameter counts. In contrast, larger models exhibit more balanced performance profiles, with GPT-4.1-mini maintaining 0.265-1.042 APD@1 across depths. Surprisingly, top performers like DeepSeek-R1 show no significant advantage on Depth4 problems (APD@1: 0.916 vs. 1.208 for DeepSeek-V3), indicating current reasoning patterns inadequately handle deep dependency chains. This suggests models could benefit from explicit dependency tracking mechanisms and iterative verification loops during code generation. For another perspective, we report APT@1 in Table 2. Notably, the trends of APT closely align with APD, reinforcing our observations of models’ depth-wise performance.
4.2 Analysis and Discussion
Dependency Structure Challenges in Multi-turn Scenarios
A deeper look of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential function compositions). However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), state-of-the-art models exhibit significant performance degradation. This phenomenon is empirically validated in Figure 5, which illustrates the Pass@1 scores across varying turn counts. The consistent performance trajectory—initially high for 1-2 turn problems followed by approximately exponential decline as turn counts increase—demonstrates the inherent challenges in multi-turn code generation. Even top-performing models (e.g., Deepseek-R1 with 20.5% Pass@1) fail to solve problems requiring more than six turns. This underscores the critical limitation to balance local correctness and global integration across iterative development cycles.
Figure 5: The Pass@1 metrics on multi-turn problems grouped by model categories and turn number.
To quantity dependency structure complexity, we introduce the Dependency Structure Complexity (DSC) metric, defined as the ratio of total turns to the maximum depth in the AST. Figure 13 presents the models’ performance across different DSC intervals, revealing that most models perform well on problems with linear dependency structures but struggle significantly as the dependency structure becomes more complex. We leave the discussion on DSC in Appendix E.1.
Fine-Grained Error Types in Multi-turn Generation
Given the significant performance gap between models in multi-turn and single-turn scenarios, we conducted studies to identify the underlying reasons. We categorized errors into three primary types: (1) Incomplete Reasoning (IR): Models often handle only straightforward "happy-path" cases and fail to generalize. They may oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapses on larger inputs. This reflects a limitation in the models’ reasoning abilities. (2) Insufficient Globalization (IG): While a function’s logic may run correctly in isolation, it may omit necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This indicates a limitation in the models’ ability to manage global context. (3) Instruction Misinterpretation (IM): Given multi-turn prompts, models could solve isolated subproblems but lack a coherent understanding of the overarching goal. Typical failures include misusing helper functions or implementing disorganized code within the top-level function, i.e. incorrect integration of components. To quantify the distribution of errors, we randomly sampled examples and manually annotated error categories. The proportion is calculated in Table 3.
Table 3: Proportion of each error type on Deepseek-V3 and o1-mini. The IR Remediation Rate is the proportion of IR errors in multi-turn scenario which are correctly answered in single-turn scenario.
Model IR(%) IG(%) IM(%) Others(%) IR Remediation Rate(%)
Deepseek-v3 47.2 14.6 32.6 5.6 16.7
o1-mini 38.7 11.8 43.0 6.5 8.3
Table 3 offers several insights. First, Incomplete Reasoning remains the dominant error type which is the main challenge of competition-level problems. To distinguish the effect of multi-turn scenario, we further introduce IR Remediation Rate to quantify how much of IR can be fixed when switching from multi-turn to single-turn scenarios. The statistics show that, most IR errors (30% of all error types) originate from the multi-turn scenario. Furthermore, Insufficient Globalization and Instruction Misinterpretation are inherently specific to multi-turn scenarios. The high rate of IM errors highlights that models struggle with consistent instruction following across multiple turns, or the codeflow task. Finally, IG errors reveal that models often lack "global awareness", highlighting the importance of both local correctness and global coherence evaluation. We leave case studies in Appendix E.2.
@@ 5 Conclusion
This paper introduces CodeFlowBench, the first benchmark specifically designed to evaluate multi-turn iterative code generation capabilities in realistic development workflows, i.e., the codeflow. Comprising 5,258 competition-level programming problems curated from Codeforces, our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APT@k, APD@k, and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 60% performance degradation). Our fine-grained analysis identifies dominant failure modes and provides insights for further advancements. Extensive experiments across 16 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also paves the way for more realistic and powerful code generation systems. Future Work: CodeFlowBench currently remains confined to competition-level problems, which should be expanded with repository-level challenges. We also plan to design and train dedicated code generators tailored to the codeflow task.
Appendix
[…]
Appendix A Features of CodeFlowBench
We collect problems from Codeforces, which offer following advantages to our benchmark:
High difficulty.
Even state-of-the-art LLMs have low pass rates on Codeforces problems [32]. Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful.
High quality.
Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.
Periodic updates.
To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly. Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.
Appendix B Data Curation Pipeline
B.1 Detail of Problem Scraping
@@
B.2 Detail of Solution Scraping
@@
B.3 Detail of Solution Code Generation
@@
B.4 Detail of Subproblems Generation
@@
A example of subproblems in shown in Figure 12.
@@
B.5 Detail of Test Cases Generation
@@
Appendix C Mathematical Expression of $PD@k$& $APD@k$
For a given problem, we define $PD@k$ as the expected maximum pass depth over $k$ independent trials of the model. Directly using only those $k$ results leads to high variance, just as with the $pass@k$ metric. By analogy to the unbiased estimator for $pass@k$, which leverages $n$ trials ($n>k$) to reduce variance, we derive a similar estimator for $PD@k$.
Let the pass depths from $n$ trials be:
$\{d_{1},d_{2},\dots,d_{n}\},$
and let
$d_{(1)}\leq d_{(2)}\leq\cdots\leq d_{(n)}$
denote these depths in ascending order. Then an unbiased estimator for $PD@k$ is
$PD@k\;=\;\sum_{j=k}^{n}d_{(j)}\,\frac{\binom{j-1}{k-1}}{\binom{n}{k}}.$ (1)
To see how this arises, consider sampling a random subset of size $k$ from the $n$ depths. Let $M\;=\;\max\{\,d_{i_{1}},d_{i_{2}},\dots,d_{i_{k}}\}$ be the maximum depth in that subset. An unbiased estimator of $M$ is $\mathbb{E}[M]$. By construct $\mathbb{E}[M]\;=\;PD@k$ ,this estimator is unbiased since its expected value exactly equals the true expected maximum depth. Consider all possible value of $M$, we have:
$\mathbb{E}[M]=\sum_{m}m\,P(M=m)=\sum_{j=k}^{n}d_{(j)}\,P\bigl{(}M=d_{(j)}\bigr{)}.$ (2)
The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with $k-1$ depths from the first $j-1$ smaller values:
$P\bigl{(}M=d_{(j)}\bigr{)}=\frac{\displaystyle\binom{j-1}{k-1}}{\displaystyle\binom{n}{k}}.$ (3)
Combining (3) and (4) immediately recovers (1).
Finally, we define
$APD@k=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}PD@k(p)$ (4)
the average $PD@k$ over the set $\mathcal{P}$ of all problems.
Appendix D Experiment
D.1 The Setting of Experiment
Due to the substantial size of the total question pool ($N=5258$), we implemented stratified sampling with proportional allocation across overall-depth categories to select 1,000 test samples, as detailed in Table 4. To validate sampling quality, we conducted a $\chi^{2}$ test comparing the overall-turn distributions between the population and sampled data (Table 5). The statistical analysis yielded a p-value of 0.1246, indicating no significant difference ($\alpha=0.05$) in distribution patterns. This confirms the representativeness of our sampling strategy and ensures the validity of subsequent analytical outcomes.
Table 4: Comparison of Overall-Depth Proportions Between the Population and the Sample
Overall-Depth # Population Pop. Proportion (%) # Sample Sam. Proportion (%)
1 1,488 28.3 283 28.3
2 2,751 52.3 523 52.3
3 870 16.5 165 16.5
4 125 2.4 24 2.4
$\geq$5 24 0.5 5 0.5
Total 5,258 100.0 1,000 100.0
Table 5: Comparison of Overall-Turn Proportions Between the Population and the Sample
Overall-Turn # Population Pop. Proportion (%) # Sample Sam. Proportion (%)
1 1,488 28.3 283 28.3
2 2,158 41.0 437 43.7
3 990 18.8 188 18.8
4 402 7.6 58 5.8
5 149 2.8 21 2.1
6 46 0.9 6 0.6
$\geq$7 25 0.5 7 0.7
Total 5,258 100.0 1,000 100.0
D.2 Instruction Templates
Multi-turn Test You are a Programming Expert. You always provide correct and reliable code solutions. You will be provided with the background of the whole problem, a programming problem and may also some pre-implemented functions.If pre-implemented functions provided, you need to call the pre-implemented functions and write a new function to solve the problem. ## Background of the whole problem: {problem_description} ## Problem Description: You need to complete name function. {statement} ## Dependency information: To solve the problem, you need to utilize the ## Pre-implemented functions {dependencies} provided. ## Pre-implemented functions: {history} ## Guidelines: - Ensure the function is executable and meets the requirement. - Handle ## Dependency information correctly. - Provide clear and concise comments to explain key parts of the code. Return your response by filling the function body following the function signature provided. Just generate the function and don’t output any examples.
In the instruction template for multi-turn testing, we implemented a set of basic heuristics to adapt to different problem types. For example, if a question has no dependencies, we omit the “## Dependency information” section. If no prior code is provided—which is common when the overall depth is 1—we exclude the “## Pre-implemented functions” section. Furthermore, if it is the final turn of the code, we append the following snippet:
import sys
def {name}():
input = sys.stdin.read().split()
[…]
Similarly, if the problem does not have any dependencies, we will also omit the section related to {dependencies}.
D.3 Pass Turn Result
Table 6: Multi-turn pass@turn results for various models.
Model Average Turn-1 Turn-2 Turn-3 Turn-4 Turn-5
Closed-Source
o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
Open-Source (7B-Level)
Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
Open-Source (32B-Level)
Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
Open-Source (70B-Level)
Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
Open-Source (Large Model)
Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
Appendix E Discussion
E.1 Discussion on Models’ DSC Performance
For a problem, the DSC metric is defined as:
[…]
Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST. Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure. Figure 13 presents the pass@1 scores of models across different DSC intervals. It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure. Only a few leading models are able to solve a limited number of problems with DSC values below 1.33. All models struggle significantly when faced with problems involving more complex structures.
E.2 Error Case Study